# Sample size and power calculations for causal inference in observational studies

### v3

@@ 3.2 Distribution of propensity scores
The propensity score is usually specified by a logistic regression model: $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}=\beta_{0}+\sum_{j=1}^{p}\beta_{j}X_{j}$. Theorem 1 motivates us to approximate the distribution of the scalar summary $W_{e}$ by a normal distribution $\mathsf{N}(\mu_{e},\sigma_{e}^{2})$, and thus approximate the propensity score distribution by a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$.
@@
To determine the two parameters $\mu_{e}$ and $\sigma_{e}$ of the logit-normal distribution, we would need to impose constraints in the form of summary information of the treatment assignment. The first constraint, which is usually available, is the proportion of the treatment group in the intended study: $r=\Pr(Z_{i}=1)$. The second constraint is a measure, denoted as $\phi$, of the similarity between the covariate distributions in the treatment and control groups, $\Pr(X\mid Z=1)$ versus $\Pr(X\mid Z=0)$. There are many choices of the specification of $\phi$; important selection criteria include: (i) there should be 1-1 map between $(r,\phi)$ and $(\mu_{e},\sigma_{e})$ so that the data generating process of the propensity score is unique, and (ii) easy to compute the propensity score distribution given $(r,\phi)$.
@@
The logit-normal distribution does not have a closed-form mean and thus it is generally difficult to solve for $(\mu_{e},\sigma_{e})$ given $(r,\phi)$. Instead, we propose to approximate the logit-normal distribution by a Beta distribution, which has a range on $[0,1]$ and is easy to compute. Applying the method in Aitchison and Shen, (1980), we can derive the “optimal” Beta function approximation of a logit-normal function by minimizing the Kullback-Leibler divergence between them. Specifically, for a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$ and its optimal approximated Beta distribution $\mathsf{Beta}(a,b)$, their parameters must satisfy the following relation: $\mu_{e}=\psi(a)-\psi(b)$ and $\sigma_{e}^{2}=\psi^{\prime}(a)+\psi^{\prime}(b)$, where $\psi$ and $\psi^{\prime}$ are the digamma and trigamma functions, respectively. Numerical examples in Supplementary Material B.2 illustrate that this approximation is accurate for a wide range of logit-normal distributions.
@@
If $e(X)\sim\mathsf{Beta}(a,b)$, then there is a closed-form relation between $a,b,r$: $r={a}/(a+b)$. Therefore, it suffices to identify the propensity score distribution if we find another easily computable relation between $a,b$ and the overlap measure $\phi$. The question now becomes how to specify such a $\phi$. Directly operating on the multivariate covariate $X$ is challenging; instead we operate on the propensity score $e(X)$, which is the coarsest balancing score (Rosenbaum and Rubin,, 1983). Let $f_{z}(u)$ be the density of the propensity score in arm $z$: $f_{z}(u)=\Pr(e(X)=u\mid Z=z)$. We propose to use the Bhattacharyya coefficient (Bhattacharyya,, 1943) to measure the covariate overlap between two groups:
$\phi\equiv\int_{0}^{1}\{f_{1}(u)f_{0}(u)\}^{1/2}\,\mathrm{d}u.$ (9)
@@
The Bhattacharyya coefficient $\phi$ satisfies $0\leq\phi\leq 1$, with $\phi=0$ if and only if the support of $f_{0}$ and $f_{1}$ does not overlap, whereas $\phi=1$ if and only if $f_{0}=f_{1}$ almost everywhere. Therefore, $\phi$ is an appropriate measure of overlap; we term it as the overlap coefficient hereafter. We choose the Bhattacharyya coefficient because it has a major computable advantage in our context, established in the following proposition (the proof is relegated to Supplementary Material A.3).
@@
For $e(X)\sim\mathsf{Beta}(a,b)$, the overlap coefficient $\phi$ has a closed-form expression as a function of $(a,b)$:
$\phi=\frac{\Gamma(a+0.5)}{{a}^{1/2}\Gamma(a)}\frac{\Gamma(b+0.5)}{{b}^{1/2}\Gamma(b)}.$ (10)
Therefore, given $(r,\phi)$, we can obtain the Beta parameters $(a,b)$ by solving for Equations $r={a}/(a+b)$ and (10). Specifically, based on the former, we can reparameterize $(a,b)$ as $a=kr$ and $b=k(1-r)$, under which Equation (10) becomes
$\phi(k)=\frac{\Gamma(kr+0.5)}{(kr)^{1/2}\Gamma(kr)}\frac{\Gamma(k(1-r)+0.5)}{\{k(1-r)\}^{1/2}\Gamma(k(1-r))}.$ (11)
We write $\phi(k)$ to highlight that, given $r$, the overlap coefficient $\phi$ is a function of the unknown $k$. The following proposition establishes that, given $r$, there is a 1-1 map between $\phi$ and $k$, and equivalently, a 1-1 map between $(r,k)$ and $(a,b)$. The proof is relegated to Supplementary Material A.4.
Proposition 2.
The function, $\phi(k)$, defined in (11), is monotonically increasing in $k$ if $k>\max\left\{(2r)^{-1},\{2(1-r)\}^{-1}\right\}$.
Proposition 2 allows us to use the bisection method to solve for $k$ efficiently. Interestingly, we found that many other specification of the overlap coefficient do not lead to a 1-1 map to the propensity score distribution. In summary, as long as users specify $(r,\phi)$, we can solve for the Beta parameters $(a,b)$ in a computationally efficient fashion, based on which we further reverse the Beta approximation of $e(X)$ to obtain the logit-normal distribution parameters $(\mu_{e},\sigma_{e})$.
To visualize the overlap coefficient, Figure 1 shows the distributions of $e(X)$ versus $1-e(X)$ corresponding to different combinations of $(r,\phi)$. In practice, we recommend users to combine substantive knowledge with Figure 1 to provide a reasonable input of $\phi$ given $r$. For example, as a generic rule of thumb for a balanced design with $r=0.5$, we recommend to code $\phi=(0.9,0.8,0.5)$ as good, medium, and poor overlap, respectively.
Figure 1: Distributions of $e(X)$ versus $1-e(X)$ corresponding to different combinations of $(r,\phi)$, where $e(X)$ follows a $\mathsf{Beta}(a,b)$ distribution with $a$ and $b$ determined by $(r,\phi)$.
@@ 3.3 Distribution of potential outcomes
For simplicity, we focus on continuous outcomes. The potential outcomes are usually specified as a linear model $Y(z)=\gamma_{z0}+X^{\prime}\gamma_{z}+\epsilon_{z}$, with $\mathbb{E}[\epsilon_{z}]=0$. Similarly as before, Theorem 1 motivates us to approximate the distribution of $X^{\prime}\gamma_{z}$ by a Gaussian random variable $W_{z}$.
@@
Recall that we approximated the distribution of the propensity score by $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ is a normal random variable. Therefore, the correlation between the potential outcomes and the propensity score is captured by the correlation between $W_{e}$ and $W_{z}$. This motivates us to decompose $W_{z}$ into two parts, one linear in $W_{e}$ and the other orthogonal to $W_{e}$. Specifically, we can decompose $W_{z}=a_{z}W_{e}+W_{e}^{\perp}$ such that $\mathrm{cov}(W_{e},W_{e}^{\perp})=0$. Then, the potential outcomes can be expressed as $Y(z)=a_{z}W_{e}+W_{e}^{\perp}+\epsilon_{z}$, where both $W_{e}^{\perp}$ and $\epsilon_{z}$ are unknown normal random variables. Because both variables are independent of $W_{e}$, it is impossible to further identify $W_{e}^{\perp}$ from $\epsilon_{z}$; thus, with slight abuse of notation, we combine these two terms and denote the sum as $\epsilon_{z}$, under which the potential outcomes are modeled as $Y(z)=a_{z}W_{e}+\epsilon_{z},\quad\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma^{2}_{z}).$
Here we have six unknown parameters in total to characterize the potential outcome distributions: $a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2}$. In sample size calculation settings, users usually have prior knowledge of the mean and variance of the observed outcome in each treatment group (e.g. available in Table 1 of a previous study in medical papers), denoted as $E_{z}\equiv\mathbb{E}(Y\mid Z=z)$ and $S_{z}^{2}\equiv\operatorname{\mathbb{V}}(Y\mid Z=z)$, for $z=0,1$. These are different from their counterparts to the potential outcomes $\operatorname{\mathbb{E}}\{Y(z)\}$ and $\operatorname{\mathbb{V}}\{Y(z)\}$. In addition to these four quantities, we would need two more constraints to identify the six parameters; it is natural to use the correlation between the covariates and the observed outcomes, i.e., $\rho_{z}\equiv\mathrm{cor}(Y,W_{e}\mid Z=z)$. We postpone the detailed discussion of $\rho_{z}$ to Section 3.4.
With the six specified values $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, one can solve for the unknown parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ by the following equations (the proof is relegated to Supplementary Material A.5)
$a_{z}$ $=\rho_{z}|S_{z}|\{\operatorname{\mathbb{V}}(W_{e}\mid Z=z)\}^{-1},\quad\mu_{z}=E_{z}-a_{z}\mathbb{E}(W_{e}\mid Z=z),\quad\sigma_{z}^{2}=(1-\rho_{z}^{2}){S_{z}^{2}}.$ (12)
The conditional mean $\mathbb{E}(W_{e}\mid Z=z)$ and variance $\operatorname{\mathbb{V}}(W_{e}\mid Z=z)$ do not have closed forms. We propose to calculate them numerically by the following formula
$\mathbb{E}(W_{e}\mid Z=z)$ $=\frac{\int_{\mathbb{R}}x\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x},$ (13)
$\operatorname{\mathbb{V}}(W_{e}\mid Z=z)$ $=\frac{\int_{\mathbb{R}}x^{2}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}-\left\{\mathbb{E}(W_{e}\mid Z=z)\right\}^{2}.$ (14)
Once the parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ are solved, we can fully determine the distribution of $Y(0)$ and $Y(1)$.
3.4 Correlation between outcomes and covariates
@@
We now interpret the seemingly opaque correlation parameter $\rho_{z}=\mathrm{cor}(Y,W_{e}\mid Z=z)$. Recall that the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, thus $W_{e}$ summarizes the covariates that are predictive of the treatment assignment. Therefore, $\rho_{z}$ measures the confounder-outcome association: as $\rho_{z}$ increases, confounding covariates exert a greater influence on the outcome between treatments. The parameter $\rho_{z}$ is intrinsically related to a standard and interpretable quantity: the covariate-outcome association, $R_{z}=\mathrm{cor}(Y,\widetilde{Y}\mid Z=z)$, where $\widetilde{Y}$ is the component of $Y$ that can be explained by the covariates. Under the linear outcome model $Y(z)=X^{\prime}\gamma_{z}+\epsilon_{z}$, we have $\widetilde{Y}=X^{\prime}\gamma_{z}$ and the square of the covariate-outcome association, $R_{z}^{2}$, is equivalent to the R-squared statistic $R_{z,Y\mid X}=\operatorname{\mathbb{V}}(X^{\prime}\gamma_{z}\mid Z=z)/\operatorname{\mathbb{V}}(Y\mid Z=z)$. Then, we can derive the following expression to show that $\rho_{z}$ is bounded by $R_{z}$:
$\rho_{z}^{2}$ $=\left[\frac{\mathrm{cov}(Y,W_{e}\mid Z=z)}{\{\mathbb{V}(W_{e}\mid Z=z)\}^{1/2}\{\mathbb{V}(\widetilde{Y}\mid Z=z)\}^{1/2}}\right]^{2}{\frac{\mathbb{V}(\widetilde{Y}\mid Z=z)}{\mathbb{V}(Y\mid Z=z)}}=\mathrm{cor}^{2}(\widetilde{Y},W_{e}\mid Z=z)R_{z}^{2}\leq R^{2}_{z}.$ (15)
@@
In randomized trials, covariates are independent of the treatment assignment by design and thus $\mathrm{cor}(\widetilde{Y},W_{e}\mid Z=z)=0$ and $\rho_{z}=0$. In observational studies, as the covariates associated with propensity scores differ more from the covariates associated with potential outcomes, $\mathrm{cor}(\widetilde{Y},W_{e}\mid Z=z)$ will get closer to 0. Equation (15) implies that users can bound $\rho_{z}^{2}$ by the $R_{z}^{2}$ by the R-squared value of a hypothetical regression of the outcome on the covariates, of which investigators often have an educated guess based on prior domain knowledge. Moreover, $\rho_{z}$ is invariant under linear transformation of the potential outcome models. Specifically, under homogeneous effects $Y(1)=Y(0)+\tau$, we always have $\rho_{1}=\rho_{0}$; under heterogeneous effects, simulations and real examples in Section 4 suggest that $\rho_{1}$ and $\rho_{0}$ are usually similar. Also, as will be shown in formula (17), the minimal sample size for $\tau^{\mbox{\tiny{ATE}}}$ depends on $\rho$ only through $\rho^{2}$. Therefore, in practice, we recommend to impose $\rho_{1}=\rho_{0}=\rho$ for simplicity, and conduct a sensitivity analysis of $|\rho|$ in the range $(0,|R_{Y|X}|)$ with a domain knowledge informed specification of $R_{Y|X}$, or set $\rho=0$ when such information is unavailable, which could lead to a mild underestimate of the sample size.

### v4

@@ 3.2 Distribution of propensity scores
This subsection shows how to, under a logistic model specification, uniquely identify the distribution of propensity scores with only two parameters: the treatment proportion $r$ and an overlap coefficient $\phi$ that measures the similarity of covariates between two groups. The key idea is to approximate the distribution of propensity scores by a computationally tractable Beta distribution, as elaborated below.
@@
Following the literature, we impose a logistic regression model for the propensity scores: $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}=\beta_{0}+\sum_{j=1}^{p}\beta_{j}X_{j}$. Theorem 1 motivates us to approximate the distribution of the scalar summary $W_{e}$ by a Normal distribution $\mathsf{N}(\mu_{e},\sigma_{e}^{2})$, and thus approximate the propensity score distribution by a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$.
@@
To determine the two parameters $\mu_{e}$ and $\sigma_{e}$ of the logit-normal distribution, we would need to impose constraints in the form of summary information about the treatment assignment. The first constraint, which is almost always available, is the proportion of the treatment group in the intended study: $r=\Pr(Z_{i}=1)$. The second constraint is a measure, denoted as $\phi$, of the similarity between the covariate distributions in the treatment and control groups, $\Pr(X\mid Z=1)$ versus $\Pr(X\mid Z=0)$. There are many choices of the specification of $\phi$; important selection criteria include: (i) there should be 1-1 map between $(r,\phi)$ and $(\mu_{e},\sigma_{e})$ so that the data generating process of the propensity score is unique, and (ii) easy to compute the propensity score distribution given $(r,\phi)$.
@@
The logit-normal distribution does not have a closed-form mean and thus it is generally difficult to solve for $(\mu_{e},\sigma_{e})$ given $(r,\phi)$. Instead, we propose to approximate the logit-normal distribution by a Beta distribution, which also has a range on $[0,1]$ and is easy to compute. Applying the method in Aitchison and Shen (1980), we can derive the “optimal” Beta distribution approximation of a logit-normal distribution by minimizing the Kullback-Leibler divergence between them. Specifically, for a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$ and its optimal Beta distribution approximation $\mathsf{Beta}(a,b)$, their parameters satisfy
$\mu_{e}=\psi(a)-\psi(b),\quad\sigma_{e}^{2}=\psi^{\prime}(a)+\psi^{\prime}(b)$ (8)
where $\psi$ and $\psi^{\prime}$ are the digamma and trigamma functions, respectively. Numerical examples in Appendix B.2 illustrate that this approximation is accurate for a wide range of logit-normal distributions.
@@
If $e(X)\sim\mathsf{Beta}(a,b)$, then there is a closed-form relation between $a,b,r$: r=a/(a+b)r={a}/(a+b). Therefore, it suffices to identify the propensity score distribution if we find another easily computable relation between $a,b$ and the overlap measure $\phi$. The question now becomes how to specify such a $\phi$. Directly operating on the multivariate covariate $X$ is challenging; instead we operate on the propensity score $e(X)$, which is the coarsest balancing score (Rosenbaum and Rubin, 1983). Let $f_{z}(u)$ be the density of $e(X)$ in arm $z$: $f_{z}(u)=\Pr(e(X)=u\mid Z=z)$. We propose to use the Bhattacharyya coefficient (Bhattacharyya, 1943) to measure the covariate overlap between two arms:
$\phi\equiv\int_{0}^{1}\sqrt{f_{1}(u)f_{0}(u)}\,\mathrm{d}u.$ (9)
@@
The Bhattacharyya coefficient $\phi$ satisfies $0\leq\phi\leq 1$, with $\phi=0$ if and only if the support of $f_{0}$ and $f_{1}$ does not overlap, whereas $\phi=1$ if and only if $f_{0}=f_{1}$ almost everywhere. Therefore, $\phi$ is an appropriate measure of overlap; we term it as the overlap coefficient hereafter. We choose the Bhattacharyya coefficient because it has a major computable advantage in our context, established in the following proposition.
@@
For $e(X)\sim\mathsf{Beta}(a,b)$ with $(a>0,b>0)$:
  1. (i)
    the overlap coefficient $\phi$ has a closed-form expression as a function of $(a,b)$:
    $\phi=\frac{\Gamma(a+0.5)}{{a}^{1/2}\Gamma(a)}\frac{\Gamma(b+0.5)}{{b}^{1/2}\Gamma(b)};$ (10)
  2. (ii)
    given $r$, $\phi$ is monotonically increasing in both $a$ and $b$;
  3. (iii)
    there is a 1-1 map (bijection) between $(r,\phi)$ and $(a,b)$.
The proof of (i)-(iii) is relegated to Appendix A.3-A.5, respectively. Proposition 1 establishes that, given $(r,\phi)$, we can obtain the Beta parameters $(a,b)$ by solving two equations: (i) r=a/(a+b)r={a}/(a+b), and (ii) Equation (10). Furthermore, we can use the bisection method to solve for $(a,b)$ efficiently, based on which we reverse the Beta approximation of $e(X)$ to obtain the logit-normal distribution parameters $(\mu_{e},\sigma_{e})$. It is worth noting that other specification of the overlap coefficient, e.g. the overlapping area between the density functions of two distributions (Inman and Bradley, 1989), usually does not lead to a closed-form expression of $\phi$ as a function of ($a,b$), which is crucial for verifying the 1-to-1 map between $(r,\phi)$ and ($a,b$).
For $e(x)\sim\mathsf{Beta}(a,b)$, we can use the Bayes theorem to show that $\Pr(e(x)\mid Z=1)\sim\mathsf{Beta}(a+1,b)$ and $\Pr(e(x)\mid Z=0)\sim\mathsf{Beta}(a,b+1)$. Figure 1 shows the overlaying distributions of $e(X)$ in the $Z=1$ and $Z=0$ groups corresponding to different combinations of $(r,\phi)$. Theoretically, the overlap coefficient $\phi$ is between 0 and 1. However, Figure 1 shows that there is little difference between the propensity score distribution if $\phi<0.8$, especially when $r<0.5$, which is the common scenario in observational studies. Further visual check (details omitted) reveals that, when $\min\{a,b\}<1$, the Beta distribution is U-shaped and the propensities are concentrated in the tail region (close to 0 or $1$), suggesting extremely poor overlap between groups. Calculations shows $a=b=1,1.5,2.5,5$ corresponding to $\phi=0.80,0.85,0.90,0.95$, respectively. Therefore, we set a simple heuristic rule of thumb for measuring the degree of overlap: $\phi<0.8,\in(0.8-0.9),\in(0.9-0.95),>0.95$ corresponds to very poor, poor, moderate, good overlap, respectively. In practice, users could compute $\phi$ from a pilot study, of which we provide a computational method in Appendix B.3. When a pilot study is not available, we recommend users to conduct sensitivity analysis on a range of $\phi$ valued based on domain knowledge.
Figure 1: Distributions of $e(X)$ of the treatment and control groups corresponding to different combinations of $(r,\phi)$, where $e(X)$ follows a $\mathsf{Beta}(a,b)$ distribution with $a$ and $b$ determined by $(r,\phi)$.
@@ 3.3 Distribution of potential outcomes
This subsection shows how to relate potential outcomes to propensity scores under mild model assumptions. The key idea is to project the outcomes into the space spanned by the scalar summary of the propensity score obtained in Section 3.2.
We assume a restricted mean model (RMM) with a homogeneous treatment effect $\tau$ for the potential outcomes
$Y(z)=f(X)+z\tau+\epsilon.$ (11)
Furthermore, we assume (i) homoscedasticity of the error, $\operatorname{\mathbb{V}}\{\epsilon\mid e(X)\}=\operatorname{\mathbb{V}}(\epsilon)$, and (ii) $f(X)$ and $W_{e}$ are jointly Normal. We do not impose additional distributional assumptions on $Y(z)$ or $\epsilon$. The homoscedasticity assumption states that the propensity score is not predictive of the error variance of potential outcomes; it is weaker than the canonical version $\operatorname{\mathbb{V}}(\epsilon|X)=\operatorname{\mathbb{V}}(\epsilon)$. The Normal assumption holds under various models for $f(X)$, most straightforwardly under a linear model $f(X)=\gamma_{0}+\sum_{j=1}^{p}\gamma_{j}X_{j}$. Specifically, with a linear $f(X)$, we can approximate the joint distribution of $W_{e}$ and $f(X)$ by a bivariate normal distribution, invoking a multivariate extension of the Lyapunov central limit theorem (Bentkus, 2005). More generally, when $f(X)$ is itself an approximately normal function of covariates–as is often the case when no single covariate dominates as specified in Theorem 1—the joint normality provides a reasonable approximation. We evaluate the empirical performance of our method under this assumption using simulations in Section 5.
@@
Under the joint normality of $f(X)$ and $W_{e}$, the correlation between the potential outcomes and the propensity score is captured by the correlation between $f(X)$ and $W_{e}$. Therefore, we decompose $f(X)$ into two parts, one collinear with $W_{e}$ and the other orthogonal to $W_{e}$: $f(X)=aW_{e}+W_{e}^{\perp}$ such that $\mathrm{cov}(W_{e},W_{e}^{\perp})=0$. Then, the potential outcomes can be expressed as $Y(z)=aW_{e}+W_{e}^{\perp}+z\tau+\epsilon$, where $W_{e}^{\perp}$ is independent of $W_{e}$ because $(W_{e},W_{e}^{\perp})$ is jointly Normal. By the definition of RMM, $\epsilon$ is also uncorrelated with any function of $X$, including $W_{e}$. Therefore, with a slight abuse of notation, we denote the compound error $W_{e}^{\perp}+\epsilon$ as $\epsilon$.
The key quantity in the asymptotic variance of the Hájek estimator, $Y(z)-\mathbb{E}\{Y(z)\}$, now reduces to $a\{W_{e}-\operatorname{\mathbb{E}}(W_{e})\}+\{\epsilon-\operatorname{\mathbb{E}}(\epsilon)\}$ for $z=0,1$, whose mean and variance are fully characterized by two unknown parameters, $a$ and $\sigma_{y}^{2}=\operatorname{\mathbb{V}}(\epsilon)$. To solve the parameters, we would need two inputs and constraints. It is natural to set the variance of potential outcomes in the corresponding group as an input, $S^{2}\equiv\operatorname{\mathbb{V}}\{Y(0)\}=\operatorname{\mathbb{V}}\{Y(1)\}$. For the other input, we use the correlation between the linear component of the propensity scores $W_{e}$—a summary of covariates—and the potential outcomes, i.e., $\rho\equiv\mathrm{cor}\{Y(1),W_{e}\}=\mathrm{cor}\{Y(0),W_{e}\}$. With the specified values $(S^{2},\rho)$, one can solve for the unknown parameters $(a,\sigma_{y}^{2})$ by the following equations (the proof is relegated to Appendix A.6)
$a$ $=\rho|S|\sigma_{e}^{-1},\quad\sigma_{y}^{2}=(1-\rho^{2}){S^{2}}.$ (12)
@@
We now interpret the seemingly opaque correlation parameter $\rho$ under homogeneous treatment effect. Recall that the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$; therefore, $W_{e}$ is a scalar summary of the covariates that are predictive of treatment assignment. Because confounders are covariates that are correlated with both treatment and outcome, we interpret $\rho$ as a measure of the confounder-outcome association: as $\rho$ increases, confounding covariates exert a greater influence on the outcome between treatments. The parameter $\rho$ is intrinsically related to a standard and interpretable quantity: the covariate-outcome association, $R=\mathrm{cor}\{Y(0),f(X)\}$, which measures the extent to which the outcome is driven by the covariates rather than noise. In fact, $R^{2}$ is equivalent to the R-squared statistic on the potential outcome: $\operatorname{\mathbb{V}}\{f(X)\}/\operatorname{\mathbb{V}}\{Y(0)\}$. Furthermore, we derive the following expression to show that $|\rho|$ is bounded by $|R|$:
$\rho^{2}$ $=\left[\frac{\mathrm{cov}\{Y(0),W_{e}\}}{\{\mathbb{V}(W_{e})\}^{1/2}[\mathbb{V}\{f(X)\}]^{1/2}}\right]^{2}{\frac{\mathbb{V}\{f(X)\}}{\mathbb{V}\{Y(0)\}}}=\mathrm{cor}^{2}\{f(X),W_{e}\}R^{2}\leq R^{2}.$ (13)
@@
In completely randomized trials, covariates are independent of the treatment assignment by design, so that $W_{e}$ reduces to a constant, and therefore $\mathrm{cor}(f(X),W_{e})=0$ and $\rho=0$. In observational studies, as the covariates associated with propensity scores overlap more with the covariates associated with potential outcomes, $\mathrm{cor}(f(X),W_{e})$ will get farther away from 0. Equation (13) implies that users can bound $\rho^{2}$ by the R-squared value encoding the proportion of the variance of potential outcomes explainable by covariates, of which investigators often have an educated guess based on prior domain knowledge. Li et al. (2018b) used the same R-squared statistic to quantify covariate-outcome association. In practice, one might estimate $|\rho|$ from a pilot study, or conduct a sensitivity analysis of $|\rho|$ in the range of $(0,|R|)$ based on domain knowledge, or simply set $\rho=0$ though it could lead to a mild underestimate of the sample size.